OpenStudy (bcwbubby):

A race horse took minutes to run miles. How long would it take the horse to run one mile

OpenStudy (rayep12):

it does not show how many min and how many miles

jabez177 (jabez177):

Um. Perhaps you could tell us how many mins and how many miles. :) You can use the Equation button below to input fractions. :) v

OpenStudy (aleah54):

it's c

OpenStudy (bcwbubby):

\[2\frac{ 2 }{ ?5 } 1\frac{ 1? }{ 2? }\] I'm new to this sorry

jabez177 (jabez177):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Aleah54 it's c \(\color{#0cbb34}{\text{End of Quote}}\) Hello and Welcome to OpenStudy! :) You seem new here and more than willing to help which is mighty fine since we need all the help we can get :) but you might want to take some time and read this site's code of conduct! http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. It's great to see another helper among us, so it would be sad if you got in trouble for giving out answers. Instead, you can try guiding the asker through their question to help them understand it. Hope you'll have a nice time here on OpenStudy, and I'll see ya around :) Have a Nice Day!

jabez177 (jabez177):

That is all okay. :)

jabez177 (jabez177):

Okay. So let me try and rewrite your question. :)

OpenStudy (bcwbubby):

ok :)

jabez177 (jabez177):

A race horse took 2 2/5 minutes to run 1 1/2 miles. How long would it take the horse to run one mile?

OpenStudy (ineedhelplz):

@BCWBubby Are there any answer choices?

OpenStudy (bcwbubby):

nope what do you guys think it is

jabez177 (jabez177):

In order to find out how long it'll take for the horse to run 1 mile. We simply divide the minutes by miles. \[2\frac{ 2 }{ 5 }\div1\frac{ 1 }{ 2 }\]That is what we are going to do. :)

OpenStudy (ineedhelplz):

2 2/5 = 12/5 1 1/2 = 3/2 (12/5) / (3/2) = x/1 cross multiply 3/2x = 12/5 -- divide by 3/2 x = (12/5) / (3/2) x = 12/5 * 2/3 x = 24/15 x = 1 9/15 which reduces to 1 3/5 minutes

jabez177 (jabez177):

@Ineedhelplz, we are supposed to work with the user, not just give out the whole steps and answer with it. Otherwise they do not learn. That is the wrong way to guide the user to the answer.

OpenStudy (ineedhelplz):

@jabez177 I gave him the steps, there is no reason in wasting the users time

OpenStudy (ineedhelplz):

You don't have to argue with me every time I try to help someone.

OpenStudy (bcwbubby):

thank you thank you thank you I thought it was something like that welp I'm off to work

jabez177 (jabez177):

You gave him the steps along with the answer. How does the user learn?

OpenStudy (ineedhelplz):

He sees how it's done

jabez177 (jabez177):

Yes but he does not learn it. That is the wrong way for the user to learn. That is why we explain every step and let them do it so they can learn how to do it themselves. In the real world, you will not be there to give him the steps and answer along with it.

OpenStudy (jtug6):

the point is, @Ineedhelplz is that you're giving answers. and that doesnt fly with OS policy. its wrong via the rules and it's not helping them. if someone gave me all the steps to solve an integral problem i wouldnt have the first clue how to actually solve it myself. it doesnt help. please don't do it again.

OpenStudy (ineedhelplz):

I understand the code of conduct, but I am not giving a direct answer @jtug6

jabez177 (jabez177):

Actually you were. You were giving the STEPS along with the ANSWER. That also counts as giving the Direct answer.

OpenStudy (jtug6):

x = 1 3/5 is an absolute DIRECT answer. youre giving all the steps and the included numerical value of what x is.

ILovePuppiesLol (ilovepuppieslol):

its alright guys, just know for future reference, work step by step and interact with them to make sure they learn better, ask them questions, just like i do in history, so like "What do you think the next step is" or "what do you think the product of this step would be" not saying what you are doing is wrong, it is just better for all of us, for the asker to learn better and for the helper to better interact with them and to improve as a teacher

OpenStudy (jtug6):

also, if moderators see your answers you will be warned/suspended ect.

OpenStudy (ineedhelplz):

Are you sure?

ILovePuppiesLol (ilovepuppieslol):

not for that, they won't

OpenStudy (ineedhelplz):

Why don't you ask one and see what they say

OpenStudy (ineedhelplz):

Because I'm sure they would have by now if everything I was doing was against the code of conduct...

ILovePuppiesLol (ilovepuppieslol):

directly stating whatever he/she is doing is against the rules and not asking nicely and explaining why won't help them to cooperate

OpenStudy (jtug6):

Yes. I've been warned multiple times as well. It's against rules. ONLY if its a direct answer, of course.

jabez177 (jabez177):

They do. They follow the Code of Conduct so whatever they see that is against the Code of Conduct, they will give a warning/suspension if you keep doing it.

ILovePuppiesLol (ilovepuppieslol):

Guys drop this topic, keep what they said in future reference to avoid consequences, if you choose to ignore, you may be punished, maybe not, its a risk you are going to take, @jabez177 and @jtug6 may stop nagging him about it, its his risk to take, if you see him do it again, just report it, don't argue

ILovePuppiesLol (ilovepuppieslol):

Because at this point, arguing gets both sides nowhere, understood?

jabez177 (jabez177):

Okay. :)

OpenStudy (jtug6):

Sure =)

OpenStudy (ineedhelplz):

...@jabez177 argues with me everytime I help someone out, and it does nothing

ILovePuppiesLol (ilovepuppieslol):

he agreed to stop @Ineedhelplz

jabez177 (jabez177):

I do not argue. You give a DA. I tell you not to. That is not arguing. :)

OpenStudy (ineedhelplz):

Ok >.>